# PDLP: A Practical First-Order Method for Large-Scale Linear Programming

### v1

@@ 4 Quickly finding approximately optimal but feasible solutions
In integer programming, it is typical for approximately optimal but feasible solutions to be acceptable. For example, branch and bound methods are commonly stopped as soon as the optimality gap falls below a predefined threshold (e.g., 1%) to ensure the user does not have to wait too long to receive a high-quality solution [69]. This pragmatic termination criteria motivates a new heuristic that we have developed that allows first-order methods to similarly find solutions with extremely small violation feasibility violations and moderate (e.g., 1%) duality gaps.
@@
    Restarted PDHG tends to converge faster for feasibility problems than optimality problems. By a feasibility problem, we mean a problem with no objective, i.e.,
[…]
@@
    $~{}\ell_{v}\leq x\leq u_{v},$
    Empirically, these problems can be orders of magnitude faster for PDHG to solve compared the original linear program (with nonzero objective).
@@
    Given a starting solution, restarted PDHG will converge to a nearby optimal solution. PDHG has the property that the distance (in terms of the PDHG norm) from the iterates to any optimal solution is nonincreasing [46]. Thus, if we warm start (7) from a primal solution obtained by applying PDHG to the original problem, we have reason to expect that the objective value will not substantially degrade, despite the fact that (7) ignores the objective function. Hence, if we have an approximately optimal solution obtained by PDHG, the result of solving (7) may retain the high quality objective and also be primal feasible to tight numerical tolerances. We can repeat this heuristic argument on the dual feasibility problem
    $\mathop{\rm maximize}_{y\in{\mathbb{R}}^{m},r\in{\mathbb{R}}^{n}}$ 00 0 (8)
@@
These two insights are employed to design Algorithm 4. In particular, Algorithm 4 strives to find a feasible solution (i.e., $10^{-8}$ maximum constraint violation for both the primal and dual) with a relative duality gap less than the user-defined tolerance $\epsilon_{\text{rel-gap}}>0$, i.e.,
|$\frac{\absolutevalue{c^{\top}x+p(y;\ell_{c},u_{c})+p(r;\ell_{v},u_{v})}}{\max\{\absolutevalue{c^{\top}x},\absolutevalue{p(y;\ell_{c},u_{c})+p(r;\ell_{v},u_{v})}\}}\leq\epsilon_{\text{rel-gap}}.$
Appendix B gives a theoretical basis for these ideas. In particular, we show that PDHG with a fixed frequency restart scheme can achieve an $O(\beta/H(A)\log(1/\epsilon))$ iteration bound on LP instances where $\beta$ is its largest singular value of the constraint matrix $A$ and $H(A)$ is its Hoffman constant [48]. In other words, the new theory shows a runtime bound that is a constant independent of the right hand side or objective vector. In contrast, [8] shows an $O(\beta/H(K)\log(1/\epsilon))$ iteration bound for restarted PDHG on general linear programs where $K$ is a matrix containing the $A$ matrix, objective vector and right hand side. Moreover, some dependence on the right hand side and/or objective vector is fundamental to restarted PDHG: indeed there are examples where $H(A)=1$ but restarted PDHG is slow to converge (see, [47, Table 1]).
@@
    Run PDLP on the primal feasibility problem, i.e., Equation 7, starting from $(x^{k},0)$ with the initial primal weight and step size equal to the current values from Step 1. Run for $k/8$ iterations or until the primal constraint violation is less than the termination tolerance. In the former case return to Step 1, continuing “normal” PDLP iterations. Otherwise, let $\tilde{x}$ be output primal solution.
@@
    Check if $(\tilde{x},\tilde{y})$ meets the termination criteria, if not return to Step 1, resuming “normal” PDLP iterations.
Note that operations that are triggered periodically based on iteration counts (such as termination checks, adaptive restarts — see the paragraph “Periodic checks of restarts and termination criteria” in Section 3) take place in Steps 1, 2, and 3 using the iteration count assoiated with the current Step.
@@ 5 Implementation
The implementation parallelizes the computation of projections, norms, distances, and matrix-vector multiplications by the constraint matrix $x\mapsto Ax$ and its transpose $y\mapsto A^{\top}y.$ Any fixed number of threads is supported. In order to parallelize these computations, it divides the data (matrices or vectors) into shards and distributes the shards among threads. Each shard is then handled independently and then combined. By default, the number of shards is taken to be four times the number of threads.
@@
6 Large-scale benchmark instances
@@
The problems were scaled such that
  1. (i)
    The objective vector has values $1$, 00 or $-1$.
  2. (ii)
    The right-hand side has values $1$, 00 or $-1$, except in the case that a constraint has two nonzero finite, nonmatching lower and upper bounds. In this case we rescale such that the maximum absolute value of the right hand side is one.
All problems (after scaling) are available for download at https://www.oliverhinder.com/large-scale-lp-problems, and some have generators available along with detailed descriptions of the problems in the repository https://github.com/ohinder/large-scale-LP-test-problems as indicated in the right column of Table 1.
@@
design-match Statistical matching with covariate balancing constraints for making causal inference on observational data. Loosely inspired by [92]. Yes
@@
heat-source-easy This problem considers a material with several randomly located heat sources. Given temperature measurements in at various parts of the material and a list of possible locations of the heat sources the aim is to recover the temperature distribution of the material. This is a PDE constrained optimization problem with linear partial differential equation constraints. Yes
heat-source-hard Same as heat-source-easy but significantly fewer measurement locations, more heat sources and more possible heat source locations. This makes the problem harder from a recover standpoint. Moreover, we find this problem is much harder for first-order methods than heat-source-easy. Yes
production-inventory A robust production-inventory problem [15]. Yes
qap-tho-150 An LP relaxation [4] of the tho150 quadratic assignment problem from QAPLIB [3]. Yes
qap-wil-100 An LP relaxation [4] of the wil100 quadratic assignment problem from QAPLIB [3]. Yes
@@
supply-chain A multicommodity flow problem motivated by optimizing the supply chain of a large retailer. Yes
@@ 7 Numerical experiments
In this section we present numerical results on a new test suite of large-scale LP instances that we have developed and released as part of this work. Our goal is to show that first-order solvers, in particular PDLP with feasibility polishing, can provide feasible solutions to tight tolerances that are close-to-optimal, in reasonable time and without requiring enormous amounts of memory. We will compare PDHG with and without feasibility polishing to three commercial state-of-the-art Gurobi solvers: a barrier (i.e., interior point) method, and primal and dual simplex methods.
Note: we do not compare against other first-order methods or GPU implementations. We do not compare against first-order methods, because in the conference version of this paper we tested against SCS, a state-of-the-art first-order method and obtained almost an order of magnitude speed-up and do not anticipate it being competitive in our tests. We do not compare against GPU implementations because the release of the PDLP C++ code in 2022 pre-dated the GPU implementation of PDLP [56]. Additionally, some of the instances that we test exceed the RAM limit of a single GPU. We expect that a GPU implementation would outperform the OR-Tools code in wall clock time for instances that fit in the GPU’s memory.
@@
We used the Pitt computing cluster [2] to perform our experiments. There were two types of machines that we used: (i) an AMD EPYC 7302 (Rome) processor with 16 cores (32 threads) and 256 GB of memory and (ii) an Intel Xeon Platinum 8352Y (Ice Lake) processor with 32 cores (64 threads) and 1 TB of memory. The high memory machine used two and a half times more computing credits per hour so to limit its use we first ran a solver on an instance on the machine with 256 GB of memory and then only switched to the 1 TB machine if insufficient memory was avaliable. We estimate that our experiments would have cost around \$20,000 to run on Google cloud (excluding the cost of building and rescaling the instances). We ran PDLP with 16 threads on the 256 GB machine and 32 threads on the 1 TB machine. For Gurobi, we left the thread parameter at its default value, which means that it will automatically use as many threads as the machine has up to the number of virtual processors avaliable in the machine [1].
@@
From Figure 1 we can see that PDLP enjoys significant speed-up from more threads on large instances. With $32$ threads PDLP has an almost 8×8$\times$8 × speed-up on the design-match instance, one of the largest in the set. This speed-up comes from parallelizing the sparse matrix-vector operation and vector-vector operations. For smaller instances, the overall speed-up can be much less, in fact, more threads can even be counterproductive [56, Table 8]. We have identified two areas to improve our multithreading performance: thread synchronization and inefficient use of memory bandwidth. Implementing ideas from the literature on multithreaded sparse matrix-vector (SpMV) operations (e.g., [22, 51]) would lead to further improvements.
@@
PDLP terminates when the maximum $\ell_{\infty}$ norms of the primal and dual residuals are less than $10^{-8}$, and the relative duality gap is less than $1\%$, i.e.,
|$\frac{\absolutevalue{c^{\top}x+p(y;\ell_{c},u_{c})+p(r;\ell_{v},u_{v})}}{\max\{\absolutevalue{c^{\top}x},\absolutevalue{p(y;\ell_{c},u_{c})+p(r;\ell_{v},u_{v})}\}}\leq 10^{-2}.$
We used the default Gurobi termination criteria, which after their own custom scaling, has an $\ell_{\infty}$ tolerance on the primal and dual residuals of $10^{-6}$, and objective gap tolerance of $10^{-8}$. Note that we do not believe that matching Gurobi’s termination criteria to ours would make a significant difference to the results. In particular, Gurobi’s barrier implementation mostly experiences OOM errors, and so is unaffected by the termination criteria; and simplex methods only find a primal and dual feasible solution at the optimum.
@@ Results
Table 3 shows the time in hours taken per instance for each method to solve the problems and whether the method required a high memory machine or ran out of memory (shading the cell red). From this table one can see that Gurobi barrier only solves three instances because it hits memory limits in all but three instances, whereas PDLP never exceeds 1 TB of memory and only in one instance requires an excess of 256 GB. Gurobi’s primal and dual simplex is competitive with PDLP in terms of memory requirements but slow: combined dual simplex only solves four instances within the time limit of 144 hours. Finally, we see that feasibility polishing is extremely effective: PDLP without polishing only solves two instances, PDLP with polishing solves eight of eleven instances. Table 4 shows that the relative duality gap achieved by PDLP with feasibility polishing is typically much better than the $1\%$ threshold that we targeted. In particular, in all cases it obtains duality gap of less than $0.4\%$ and other than the two shipping instances the duality gap is always less than $0.1\%$.
Table 3 shows both PDLP and Gurobi had errors. Both of Gurobi’s errors gave the error code 10025. We believe this is a bug in Gurobi as the Gurobi error codes [42] state that the ‘user has called a query routine on a model with more than 2 billion non-zero entries’ even though we did not call a query routine. We have made Gurobi aware of this issue. Both of PDLP’s errors were numerical errors that occurred within a few hundred iterations.
@@
design-match 68.0 9.3 OOM 75.3 32.4
tsp-gaia-10m TIME 3.0 28.1 90.1 40.2
tsp-gaia-100m TIME 21.1 OOM OOM OOM
heat-source-easy TIME 60.0 OOM ERR TIME
heat-source-hard TIME TIME OOM ERR TIME
production-inventory TIME TIME 5.8 TIME 43.7
qap-tho-150 ERR ERR OOM TIME TIME
qap-wil-100 44.4 0.3 OOM TIME TIME
world-shipping TIME 53.8 OOM TIME TIME
mediterranean-shipping TIME 32.6 OOM TIME TIME
@@
Table 3: Comparision of solvers in terms of the number of hours to solve instances (excluding time to read the instance). Termination criteria for Gurobi was set as default; termination criteria for PDLP was $10^{-8}$ maximum violation on primal and dual feasibility but $10^{-2}$ on the relative duality gap. Cells shaded red indicate runs that required the 1 TB RAM machine. TIME indicates that the solver exceeded 144 hours. OOM stands for ‘out of memory’, meaning the solver exceeded the 1 TB RAM limit. ERR denotes that there was an error during the solve.
@@
world-shipping $-1.18\times 10^{5}$ $3.7\times 10^{2}$ $0.16$
mediterranean-shipping $-1.07\times 10^{3}$ $6.95\times 10^{0}$ $0.33$
@@
Table 4: Objectives and duality gaps produced from PDLP with feasibility polishing for instances that successfully terminated.

### v2

@@ 4 Quickly finding approximately optimal but feasible solutions
In integer programming, approximately optimal but feasible solutions are typically acceptable. For example, branch-and-bound methods are commonly stopped as soon as the optimality gap falls below a predefined threshold (e.g., 1%) to avoid excessive computation time while still delivering high-quality solutions [76]. This pragmatic termination criterion motivates a new heuristic we develop, called feasibility polishing, that enables first-order methods to similarly find solutions with extremely small feasibility violations and moderate (e.g., 1%) duality gaps.
@@
    Restarted PDHG tends to converge faster for feasibility problems than for optimality problems. By a feasibility problem, we mean a problem with no objective, i.e.,
    $\mathop{\rm minimize}_{x\in{\mathbb{R}}^{n}}$ $~0$ (6)
@@
    $~\ell_{v}\leq x\leq u_{v}.$
    Empirically, these problems can be orders of magnitude faster for PDHG to solve compared to the original linear program (with nonzero objective).
@@
    Given a starting solution, restarted PDHG converges to a nearby optimal solution. PDHG has the property that the distance (in terms of the PDHG norm) from the iterates to any optimal solution is nonincreasing [53]. The same property also holds for restarted PDHG (e.g., see Proposition 9 of Applegate et al. [10]). Consequently, one can show, if we warm-start restarted PDHG on (6) from a primal solution obtained by applying PDLP to the original problem (and set the starting dual iterate to zero, which is trivially dual feasible for (6)), then restarted PDHG will find a nearby primal feasible solution.
    Thus, if we start from a near-optimal solution (i.e., from PDLP) as a warm-start for restarted PDHG on the primal feasibility problem (6), then the result will be a primal solution obtaining a similar objective value which satisfies tight feasibility tolerances. This idea is visualized in Figure 1.
Figure 1: Visualization of feasibility polishing.
We can repeat this heuristic argument on the dual feasibility problem
$\mathop{\rm maximize}_{y\in{\mathbb{R}}^{m},r\in{\mathbb{R}}^{n}}$ $0$ (7)
@@
Appendix C gives a theoretical basis for these two ideas. In particular, we show that PDHG with a fixed-frequency restart scheme and fixed step size can achieve an $O(\beta/H(A)\log(1/\epsilon))$ iteration bound on LP instances, where $\beta$ is the largest singular value of the constraint matrix $A$ and $H(A)$ is its Hoffman constant [55]. In other words, the new theory shows a runtime bound that is constant and independent of the right-hand side or objective vector. In contrast, [10] shows an $O(\beta/H(K)\log(1/\epsilon))$ iteration bound for restarted PDHG on general linear programs, where $K$ is a matrix containing the $A$ matrix, objective vector, and right-hand side. Moreover, some dependence on the right-hand side and/or objective vector is fundamental to restarted PDHG: indeed, there are examples where $H(A)=1$ but restarted PDHG is slow to converge (see [54, Table 1]). We also show that our polishing algorithm will converge to a nearby feasible solution.
@@
These two insights are employed to design Algorithm 4. In particular, Algorithm 4 strives for a solution that satisfies extremely tight feasibility tolerances (e.g., $10^{-8}$ relative tolerances) but with a moderately large user-defined tolerance $\epsilon_{\text{rel-gap}}>0$ on the relative duality gap, i.e.,
[…]
e.g., with $\epsilon_{\text{rel-gap}}=10^{-2}$. While this is a somewhat nonstandard relative gap definition (i.e., more typical would be a max over the two terms in the denominator), we use this choice for legacy reasons.
@@
    Run PDLP on the primal feasibility problem, i.e., Equation 6, starting from $(x^{k},0)$ with the initial primal weight and step size equal to the current values from Step 1. Run for $k/8$ iterations or until the primal constraint violation is less than the termination tolerance. In the former case, return to Step 1, continuing “normal” PDLP iterations. Otherwise, let $\tilde{x}$ be the output primal solution.
@@
[…]
Note that operations that are triggered periodically based on iteration counts (such as termination checks, adaptive restarts — see the paragraph “Periodic checks of restarts and termination criteria” in Section 3) take place in Steps 1, 2, and 3 using the iteration count associated with the current Step.
@@ 5 Implementation
The implementation parallelizes the computation of projections, norms, distances, and sparse matrix-dense vector multiplications by the constraint matrix $x\mapsto Ax$ and its transpose $y\mapsto A^{\top}y.$ Any fixed number of threads is supported. To parallelize dense vector and vector-vector operations, it divides the vector into (approximately) equal sized shards, and performs each shard’s operations independently in parallel (using Eigen for the per-shard operation). To parallelize sparse matrix-dense vector multiplications, it divides the sparse matrix by rows into shards with (approximately) the same number of nonzeros, and performs each shard’s product independently in parallel (using Eigen’s sparse matrix-dense vector product for the per-shard operation). Each shard is then handled independently and then combined [^5] The algorithm is deterministic given a fixed number of shards. In each threaded operation, the per-shard operations are fully independent, so execution order doesn’t matter, and the per-shard results are processed sequentially (and deterministically). Changing the number of shards changes the order of the floating-point operations, and can result in different round-off errors.. By default, the number of shards is taken to be four times the number of threads.
@@
6 A New Collection of Large-Scale Instances
@@
design-match Randomly generated statistical matching problem with covariate balancing constraints for making causal inference on observational data. Loosely inspired by Zubizarreta [100]. Yes
@@
heat-source-easy This problem considers a material with several randomly located heat sources. Given temperature measurements at various parts of the material and a list of possible locations of the heat sources, the aim is to recover the temperature distribution of the material. This is a PDE constrained optimization problem with linear partial differential equation constraints. Yes
heat-source-hard Same as heat-source-easy but significantly fewer measurement locations, more heat sources and more possible heat source locations. This makes the problem harder from a recovery standpoint. Moreover, we find this problem is much harder for first-order methods than heat-source-easy. Yes
production-inventory A robust production-inventory problem [16]. Data is randomly generated. Yes
qap-tho-150 An LP relaxation [4] of the tho150 quadratic assignment problem from QAPLIB [3]. Data is randomly generated. Yes
qap-wil-100 An LP relaxation [4] of the wil100 quadratic assignment problem from QAPLIB [3]. Data is randomly generated. Yes
@@
supply-chain A multicommodity flow problem motivated by optimizing the supply chain of a large retailer. Data is randomly generated. Yes
@@
All these instances were rescaled such that
  1. (i)
    The objective vector has values 1-1, 0, or $1$.
  2. (ii)
    The right-hand side has values -$\infty$, 1-1, 0, 11, or $\infty$, except in the case that a constraint has two nonzero, finite, nonmatching lower and upper bounds. In this case, we rescale such that the maximum absolute value of the right hand side is one.
These instances (after scaling) are available for download at
https://www.oliverhinder.com/large-scale-lp-problems
and some have generators available along with detailed descriptions of the problems in the repository
https://github.com/ohinder/large-scale-LP-test-problems
as indicated in the right column of Table 1. Although, all the instances in this repository are randomly generated (i.e., those marked with full details available), great effort was made by the authors to make this generation as realistic as possible, coupled with clearly documented reasoning.
@@ 7 Numerical experiments
In this section we present numerical results on a new test suite of large-scale LP instances that we have developed and released as part of this work. Our goal is to show that first-order solvers, in particular PDLP with feasibility polishing, can provide feasible solutions to tight tolerances that are close-to-optimal, in reasonable time and without requiring enormous amounts of memory. We will compare PDLP with and without feasibility polishing to three commercial state-of-the-art Gurobi solvers: a barrier (i.e., interior point) method, and primal and dual simplex methods.
We also provide some supplementary experiments in Appendix F on smaller test problems. These experiments (i) demonstrate the value of feasibility polishing on the Mittelmann benchmark and report the performance of PDLP in detecting infeasibility on a standard collection of primal infeasible problems. Not surprisingly we find that Gurobi is much faster than PDLP on most of these relatively small instances.
Detailed CSVs of the results and log files for all experiments can also be found at:
$\text{\url{https://www.oliverhinder.com/large-scale-lp-problems}}.$
We use the PDLP implementation in OR-Tools version 9.10. Instructions on how to run PDLP to reproduce our results can be found at
https://github.com/ohinder/large-scale-LP-test-problems/blob/main/SOLVER-INSTRUCTIONS.md.
Note: we do not compare against other first-order methods or GPU implementations. We do not compare against first-order methods, because in the conference version of this paper [9] we already tested against SCS, a state-of-the-art first-order method. In these tests, PDLP was $19\times$ faster on NETLIB [9, Table 24] and $6.3\times$ faster on MIPLIB 2017 relaxations, and thus we do not anticipate it being competitive in our tests. We do not compare against GPU implementations because the release of the PDLP C++ code in 2022 pre-dated the GPU implementation of PDLP [66]. Additionally, some of the instances that we test exceed the RAM limit of a single GPU. We expect that a GPU implementation would outperform the OR-Tools code in wall clock time for instances that fit in the GPU’s memory.
@@
We used the Pitt computing cluster [2] to perform our experiments. There were two types of machines that we used: (i) an AMD EPYC 7302 (Rome) processor with 16 cores (32 threads) and 256 GB of memory and (ii) an Intel Xeon Platinum 8352Y (Ice Lake) processor with 32 cores (64 threads) and 1 TB of memory. The high memory machine used two and a half times more computing credits per hour so to limit its use we first ran a solver on an instance on the machine with 256 GB of memory and then only switched to the 1 TB machine if insufficient memory was available. We estimate that our experiments would have cost around \$20,000 to run on Google cloud (excluding the cost of building and rescaling the instances). We ran PDLP with 16 threads on the 256 GB machine and 32 threads on the 1 TB machine. For Gurobi, we left the thread parameter at its default value, which means that it will automatically use as many threads as the machine has up to the number of virtual processors available in the machine [1].
@@
From Figure 2 we can see that PDLP enjoys significant speed-up from more threads on large instances. With $32$ threads PDLP has an almost $8\times$ speed-up on the design-match instance, one of the largest in the set. This speed-up comes from parallelizing the sparse matrix-vector operation and vector-vector operations. For smaller instances, the overall speed-up can be much less, in fact, more threads can even be counterproductive [66, Table 8]. We have identified two areas to improve our multithreading performance: thread synchronization and inefficient use of memory bandwidth. Implementing ideas from the literature on multithreaded sparse matrix-vector (SpMV) operations (e.g., [24, 58]) would lead to further improvements.
@@
To test feasibility polishing, our experiments do not use the default PDLP termination criteria. Instead, we use a tighter criterion for feasibility and a looser criterion for the optimality gap. In particular, in our experiments PDLP terminates when the maximum $\ell_{\infty}$ norms of the primal and dual residuals are less than $10^{-8}$, i.e., $x\in\mathcal{X}$, $y\in\mathcal{Y}$, $r\in\mathcal{R}$,
$\ell_{c}-10^{-8}\leq Ax\leq u_{c}+10^{-8},\quad\|c-A^{\top}y-r\|_{\infty}\leq 10^{-8},$ (9)
and the relative duality gap is less than $1\%$, i.e.,
[…]
Recall, however, that all linear programs have been scaled as per Section 6 which impacts the termination criteria. Specifically, Equation 9 translated to the original linear programs becomes equivalent to
[…]
and
[…]
Finally, we include the infeasibility termination criteria in Appendix D.
We used the default Gurobi termination criteria, which after presolve and their own custom scaling, has an $\ell_{\infty}$ tolerance on the primal and dual residuals of $10^{-6}$, and objective gap tolerance of $10^{-8}$. Note that we do not believe that making Gurobi’s termination criteria match ours would make a significant difference to the results. In particular, Gurobi’s barrier implementation mostly experiences OOM errors, and so is unaffected by the termination criteria; and simplex methods only find a primal and dual feasible solution at the optimum.
@@ Results
Table 3 shows the time in hours taken per instance for each method to solve the problems and whether the method required a high memory machine or ran out of memory (shading the cell red). From this table one can see that Gurobi barrier only solves three instances because it hits memory limits in all but three instances, whereas PDLP never exceeds 1 TB of memory and only in one instance requires an excess of 256 GB. Gurobi’s primal and dual simplex is competitive with PDLP in terms of memory requirements but slow: combined primal and dual simplex only solves four instances within the time limit of 144 hours. Finally, we see that feasibility polishing is extremely effective: PDLP without polishing only solves two instances, PDLP with polishing solves eight of eleven instances. Table 4 shows that the relative duality gap achieved by PDLP with feasibility polishing is typically much better than the $1\%$ threshold that we targeted. In particular, in all cases it obtains a duality gap of less than $0.4\%$ and other than the two shipping instances the duality gap is always less than $0.1\%$. Of course, these duality gaps are much worse than what Gurobi would achieve if it solved the problem (recall $10^{-8}$ is the default absolute optimality gap for which Gurobi barrier terminates) but for many applications this may be satisfactory.
Table 3 shows that both PDLP and Gurobi had errors. Both of Gurobi’s errors gave the error code 10025. We believe this is a bug in Gurobi as the Gurobi error codes [49] state that the ‘user has called a query routine on a model with more than 2 billion non-zero entries’ even though we did not call a query routine. We have made Gurobi aware of this issue. Both of PDLP’s errors were numerical errors that occurred within a few hundred iterations.
@@
[…]
tsp-gaia-10m TIME $\cellcolorred!30$28.1 90.1 40.2
[…]
qap-wil-100 44.4 $\cellcolorred!30$OOM TIME TIME
[…]
@@
Table 3: Comparison of solvers in terms of the number of hours to solve instances (excluding time to read the instance). Termination criteria for Gurobi was set as default; termination criteria for PDLP was $10^{-8}$ maximum violation on primal and dual feasibility but $10^{-2}$ on the relative duality gap. Cells shaded red indicate runs that required the 1 TB RAM machine. TIME indicates that the solver exceeded 144 hours. OOM stands for ‘out of memory’, meaning the solver exceeded the 1 TB RAM limit. ERR denotes that there was an error during the solve.
@@
world-shipping $1.18\times 10^{5}$ $3.7\times 10^{2}$ $0.16$
mediterranean-shipping $1.07\times 10^{3}$ $6.95\times 10^{0}$ $0.33$
@@
Table 4: Objectives and duality gaps produced from PDLP with feasibility polishing for instances that successfully terminated. Note the relative gap is computed with the somewhat non-standard formula given in Equation 8.
8 Conclusion
In this work, we demonstrated that feasibility polishing significantly improves the ability of PDLP to find solutions with tight feasibility tolerances (both primal and dual) but more relaxed optimality tolerances. Practitioners would obviously prefer solutions that satisfy tight feasibility and optimality tolerances—solutions which are reliably provided by interior point methods and the simplex method on moderately large instances. However, the standard strategy of relaxing both feasibility and optimality tolerances to allow first-order methods to produce solutions within a reasonable timeframe [75] is undesirable: moderate feasibility violations can give unexpectedly catastrophic solutions to the true underlying problem. This issue is not merely hypothetical; it occurs in the ‘irish-electricity’ instance as discussed in the introduction. Moreover, ruling out this issue requires a solution with tight feasibility tolerances, which defeats the purpose of using a first-order method for a fast, inaccurate solve.
We also showed, on a new collection of very large linear programs and a machine with 1TB of RAM, that PDLP with feasibility polishing can find solutions with extremely small feasibility violations and moderate duality gaps for several problems that Gurobi is unable to solve within six days. While we expect our work to expand the number of large-scale linear programs that can be solved, it is important to emphasize the inherent limitations of first-order methods: they are unlikely to ever completely replace second-order methods. In particular, one can construct linear systems with binary entries and condition number exponential in the dimension [6]. Such linear systems are unlikely to be solved efficiently by first-order methods, which in the worst case require a number of iterations polynomial in the condition number to obtain high-accuracy solutions.
Nonetheless, our numerical results illustrate that large sets of practical linear programs can be effectively tackled with first-order methods. Indeed, when constraint matrices are totally unimodular [54] or have small circuit imbalance [29], first-order methods provably require only a polynomial (in the problem dimensions) number of iterations to find solutions satisfying the optimality conditions with tight tolerances. Relaxing the termination criteria to require only tight feasibility tolerances with moderate duality gap tolerances will likely expand this set even further.